did I work this problem correctly
3(3x-1)=3x-2(3-x)
9x - 3= 3x - 6 + 2x
9x -3 =5x - 6
4x-3=-6
x=-.75

Answer by bucky(2100)

3(3x-1)=3x-2(3-x)
9x - 3= 3x - 6 + 2x <----OK
9x -3 =5x - 6 <---- OK
4x-3=-6 <---- OK
.
Let me see. Add 3 to both sides to get:
.
4x = -3
.
Then divide both sides by 4 and you get:
.
x=-.75 <---- Yes! Good Job! Keep up the good work.